OpenStudy (anonymous):

find domain: y=x-6/x^2-5x-36

OpenStudy (luigi0210):

First what you need to do is factor the numerator and denominator.

OpenStudy (anonymous):

how?

OpenStudy (luigi0210):

To see if they have any common factors :P

OpenStudy (anonymous):

\[y=\frac{x-6}{x^2-5x-36}=\frac{x-6}{(x-9)(x+4)}\] Any value of \(x\) would work, except for those that make the denominator 0.

OpenStudy (anonymous):

yea um just hopping in this lesson can u teach me like by steps

OpenStudy (anonymous):

ok got it so the domain would be 9 and -4?

OpenStudy (anonymous):

sorry x would not be 9 or -4

OpenStudy (anonymous):

Yes. So you would say something like "all real numbers \(x\) except for 9 and -4."

OpenStudy (anonymous):

ok thanks. so what if the problem has a radical in it like √3x-10/ Is the answer x is > or = 10/3?

OpenStudy (anonymous):

SithsAndGiggles? Still There?

OpenStudy (anonymous):

If \(f(x)=\sqrt{3x-10}\), then for \(f\) to be defined, you have the restriction that \(3x-10\ge0\), or \(x\ge\dfrac{10}{3}\). This would then be your domain, so yes you're right.